OpenStudy (andrewkaiser333):

The line through (-4,2) with the slope of 3/4 in point-slope form.

OpenStudy (andrewkaiser333):

please help

OpenStudy (andrewkaiser333):

i was sick and don't remember how to do this

OpenStudy (andrewkaiser333):

i want the who thing so i know how ok?

OpenStudy (anonymous):

slope is m =3/4 and (-4,2) is x and y coordinates substitute them in y=mx +C and you find C then you need to rewrite the m and C in the y=mx+C so this is how it goes: y=2 x=-4 and m = 3/4 so 2=(-4x3/4) +C then 2=-4 +C then 2+4 = C so C=6 Now substitute m=3/4 and C=6 in the equation y=(3/4)x+6 now multiply both sides by 4 to make it look better 4y=3x+24

OpenStudy (andrewkaiser333):

ph ok thanks

OpenStudy (andrewkaiser333):

m=5 right?

OpenStudy (anonymous):

m is given as 3/4 it is not 5 I don't know what you are saying